OpenStudy (anonymous):

solutions of 2^x +3^x +4^x -5^x=0???

OpenStudy (anonymous):

Are you misreading, Anusha? The equation is: \[2^x+3^x+4^x-5^x=0\]

OpenStudy (anonymous):

Consider numerical testing. Test the equation at \(x=1\). Now test it at \(x=2\). Then try \(x=3\). See the idea?

OpenStudy (amistre64):

log(0) does not exist; so logging both sides is futile

OpenStudy (anonymous):

Do you want integral solutions?

OpenStudy (shubhamsrg):

and you'll require the knowledge of high degree interpolation i guess in this type of ques..

OpenStudy (anonymous):

Oh my good lord, this is the root of a ridiculous exponential equation.

OpenStudy (anonymous):

You will definitely need interpolation.

OpenStudy (anonymous):

ya the solution can be of any type it can be integral or complex

OpenStudy (anonymous):

interpolation???

OpenStudy (anonymous):

Interpolation is how you find the approximation of a solution to an equation by narrowing the interval of potential solutions over and over.

OpenStudy (anonymous):

hmm thanks for the help:)

OpenStudy (amistre64):

i always called it "trial and error"

OpenStudy (anonymous):

Hmmmm. @Limitless What if I take all the terms(2^x, 3^x....) in terms of a common factor and then solve it.?

OpenStudy (anonymous):

It is definitely not looking pretty in any way what-so-ever. I don't think it was intended to be solved in an analytical way. Maybe I am wrong.

OpenStudy (amistre64):

i see one place we can simplify 2^x +3^x +4^x -5^x=0 ; 4^x = 2^(2x) 2^x +3^x + 2^(2x) -5^x=0 but im not sure of thats useful or not :/

OpenStudy (amistre64):

id say give it a shot, the worst that could happen is it doesnt pan out; and even if its not the right method, you might even stumble upon a right answer :)

OpenStudy (amistre64):

its all in the dbl chk

OpenStudy (anonymous):

Okay. Damn. I'm sorry. It doesnt turn into a linear equation. My bad.

OpenStudy (amistre64):

2^x +3^x +4^x -5^x=0 2^x +2^(2x) = 5^x - 3^x 2^x( 1 + 2^x) = 5^x - 3^x 1 + 2^x = (5^x - 3^x)/2^x http://www.wolframalpha.com/input/?i=2%5Ex+%2B3%5Ex+%2B4%5Ex+-5%5Ex%3D0 ewww ....

OpenStudy (anonymous):

Well.. There is an exponential equation of which it is a root. I thought you might be onto something, but I could not tell.

OpenStudy (amistre64):

can we turn it into a taylor and give it a shot?

OpenStudy (anonymous):

I'm not sure how we'd convert. I tried using the binomial theorem with every term in terms of \(2^x\). It does not look pretty nor useful. Also, I can't remember if there is an \(n\)th derivative of the general exponential function..

OpenStudy (amistre64):

a^x (ln(a))^n i think

OpenStudy (anonymous):

Then there is certainly a taylor series. It would be simply \[\sum_{n=0}^{\infty}\frac{f^{(n)}(a)}{n!}(x-a)^n=0\] where we set \(f(x)=2^x+3^x+4^x-5^x\). I am not entirely sure what would be the best choice of \(a\).

OpenStudy (amistre64):

a = 0 or 1 is my idea; but im sure there is an interval of convergence to consider

OpenStudy (anonymous):

I don't know what the interval of convergence is for a general taylor series off of the top of my head. Would it be just like the radius of convergence for a power series?

OpenStudy (amistre64):

yes, since a taylor is just a powr series

OpenStudy (amistre64):

ln2 + ln3 + ln4 - ln5 = ln(2.3.4/5) = ln(24/5) if we pick a=0

OpenStudy (anonymous):

I'm trying to think we need to consider the interval of convergence first. I don't know how, though.. >_<

OpenStudy (anonymous):

The fact that \(a\) is part of the coefficient has totally thrown me off.

OpenStudy (amistre64):

\[T(x) = 2 + \sum_{n=1}^{inf}\frac{ln^n(24/5)x^n}{n!}\] \[2+\lim\frac{ln^n(24/5)x^n}{n!}*\frac{(n-1)!}{ln^{(n-1)}(24/5)x^{(n-1)}}\] \[2+\lim\frac{ln(24/5)x}{n}\] \[2+|ln(24/5)x|\lim\frac{1}{n}\] as n to inf; thats 0 and 0<1 always so the interval might be everywhere convergent ..... just a guess tho

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=sum%28n+from+1+to+inf%29+%28ln%2824%2F5%29%29%5Enx%5En%2Fn%21+and+2%5Ex%2B3%5Ex%2B4%5Ex-5%5Ex-2 its not everywhere convergent sooo, i missed it in the figuring :)

OpenStudy (anonymous):

It seems to be badly approximating our function, too... http://www.wolframalpha.com/input/?i=sum%28n+from+1+to+inf%29+%28ln%2824%2F5%29%29%5Enx%5En%2Fn%21+and+2%5Ex%2B3%5Ex%2B4%5Ex-5%5Ex